# Simulating Polynomial-Time Nondeterministic Turing Machines via Nondeterministic Turing Machines

### v22

@@ Abstract.
We prove in this paper that there is a language $L_{s}$ accepted by some nondeterministic Turing machine that runs within time $O(n^{k})$ for any positive integer $k\in\mathbb{N}_{1}$ but not by any ${\rm co}\mathcal{NP}$ machines. Then we further show that $L_{s}$ is in $\mathcal{NP}$, thus proving that
@@ 1.1. Our Contributions
$\mathcal{P}\neq\mathcal{NP}$. [^1] So, from this paper, we know that resolving the $\mathcal{P}\overset{?}{=}\mathcal{NP}$ problem does not necessarily require diagonalization techniques.
@@ 1.2. Background and Prior Work
In history, the fundamental measure of time opened the door to the study of the extremely expressive time complexity class $\mathcal{NP}$, one of the most important classical complexity classes, i.e., nondeterministic polynomial-time. Specifically, $\mathcal{NP}$ is the set of decision problems for which the problem instances, where the answer is “yes,” have proofs verifiable in polynomial time by some deterministic Turing machine, or alternatively the set of problems that can be solved in polynomial time by a nondeterministic Turing machine (see e.g., [A2]). The famous Cook-Levin theorem [Coo71, Lev73] shows that this class has complete problems, which states that Satisfiability is $\mathcal{NP}$-complete, i.e., Satisfiability is in $\mathcal{NP}$ and any other language in $\mathcal{NP}$ can be reduced to it in polynomial time. At the same time, it is also worth noting that this famous and seminal result opened the door to research into the renowned rich theory of $\mathcal{NP}$-completeness [Kar72].
@@
In the next few paragraphs, let us recall some of the interesting background, the current status, the history, and the main goals of the field of proof complexity. On the one hand, as we all know well, proof theory is a major branch of mathematical logic and theoretical computer science in which proofs are treated as formal mathematical objects, facilitating their analysis by mathematical techniques (see e.g., [A4]). Some of the major areas of proof theory include structural proof theory, ordinal analysis, automated theorem proving, and proof complexity, and so forth. It should be pointed out that much research also focuses on applications in computer science, linguistics, and philosophy, from which we see that the impact of proof theory is very profound. Moreover, as pointed out in [Raz04] by Razborov, proof and computations are among the most fundamental concepts relevant to virtually any human intellectual activity. Both have been central to the development of mathematics for thousands of years. The effort to study these concepts themselves in a rigorous, metamathematical way was initiated in the $20$th century and led to the flourishing of mathematical logic and derived disciplines.
@@
On the other hand, according to the points of view given in [Rec76], logicians have proposed a great number of systems for proving theorems, which give certain rules for constructing proofs and for associating a theorem (formula) with each proof. More importantly, these rules are much simpler to understand than the theorems. Thus, a proof gives a constructive way of understanding that a theorem is true. Specifically, a proof system is sound if every theorem is true, and it is complete if every true statement (from a certain class) is a theorem (i.e., has a proof); see e.g., [Rec76].
@@
Indeed, the relationship between proof theory and proof complexity is inclusion. Within the area of proof theory, proof complexity is the subfield aiming to understand and analyze the computational resources that are required to prove or refute statements (see e.g., [A5]). Research in proof complexity is predominantly concerned with proving proof-length lower and upper bounds in various propositional proof systems. For example, among the major challenges of proof complexity is showing that the Frege system, the usual propositional calculus, does not admit polynomial-size proofs of all tautologies; see e.g., [A5]. Here the size of the proof is simply the number of symbols in it, and a proof is said to be of polynomial size if it is polynomial in the size of the tautology it proves. Moreover, contemporary proof complexity research draws ideas and methods from many areas in computational complexity, algorithms, and mathematics. Since many important algorithms and algorithmic techniques can be cast as proof search algorithms for certain proof systems, proving lower bounds on proof sizes in these systems implies run-time lower bounds on the corresponding algorithms; see e.g., [A5].
@@
Likewise, propositional proof complexity is an area of study that has seen a rapid development over the last two decades, which plays as important a role in the theory of feasible proofs as the role played by the complexity of Boolean circuits in the theory of efficient computations; see, e.g., the celebrated work [Raz15] by Razborov. In most cases, according to [Raz15], the basic question of propositional proof complexity can be reduced to that given a mathematical statement encoded as a propositional tautology $\psi$ and a class of admissible mathematical proofs formalized as a propositional proof system $P$, what is the minimal possible complexity of a $P$-proof of $\psi$? In other words, propositional proof complexity aims to understand and analyze the computational resources required to prove propositional tautologies, in the same way that circuit complexity studies the resources required to compute Boolean functions. In light of the point of view given in [Raz15], the task of proving lower bounds for strong proof systems like Frege or Extended Frege modulo any hardness assumption in the purely computational world may be almost as interesting; see e.g., [Raz15]. It is worth noting that it has been observed by Razborov [Raz15] that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$ implies lower bounds for any propositional proof system. In addition, see, e.g., [Raz03] for interesting directions in propositional proof complexity.
@@
Let us return back to the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem for this moment. As a matter of fact, the so-called $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem is the central problem in proof complexity; see, e.g., the excellent survey [Kra19]. It formalizes the question of whether or not there are efficient ways to prove the negative cases in ours and in many other similar examples. Moreover, the ultimate goal of proof complexity is to show that there is no universal propositional proof system allowing for efficient proofs of all tautologies, which is equivalent to showing that the computational complexity class $\mathcal{NP}$ is not closed under the complement; see e.g., [Kra19].
Particularly, in 1974, Cook and Reckhow showed in [CR74] that there exists a super proof system if and only if $\mathcal{NP}$ is closed under complement; that is, if and only if $\mathcal{NP}={\rm co}\mathcal{NP}$. It is, of course, open whether any superproof system exists. The above result has led to what is sometimes called “Cook’s program” for proving $\mathcal{NP}\neq{\rm co}\mathcal{NP}$: prove superpolynomial lower bounds for proof lengths in stronger and stronger proposition proof systems until they are established for all abstract proof systems. Cook’s program, which also can be seen as a program to prove $\mathcal{P}\neq\mathcal{NP}$, is an attractive and plausible approach; unfortunately, it has turned out to be quite hard to establish superpolynomial lower bounds on common proof systems; see, e.g., the excellent survey on proof complexity [Bus12].
@@ 2. Preliminaries
There exists a Turing machine $U$ such that for every $x,\alpha\in\{0,1\}^{*}$, $U(x,\alpha)=M_{\alpha}(x)$, where $M_{\alpha}$ denotes the Turing machine represented by $\alpha$. Moreover, if $M_{\alpha}$ halts on input $x$ within $T(|x|)$ steps, then $U(x,\alpha)$ halts within $cT(|x|)\log T(|x|)$  [^5] In this paper, $\log n$ stands for $\log_{2}n$. steps, where $c$ is a constant independent of $|x|$ and depending only on $M_{\alpha}$’s alphabet size, number of tapes, and number of states.
@@ 3. All Polynomial-time Nondeterministic Turing Machines Are Enumerable
There is another way to enumerate all polynomial-time nondeterministic Turing machines without encoding the order of polynomial into their representation. To do so, we need the Cantor pairing function (see Fig. 2 below from [A8]):
@@ 4. Simulation of All Polynomial-time Nondeterministic Turing Machines
One may wish to apply simulation techniques to the case of $\mathcal{P}$ versus ${\rm co}\mathcal{P}$ and doubt that simulation techniques are unable to produce a language not in ${\rm co}\mathcal{P}$. In this scenario, we remark that the simulation technique is unable to create a language not in ${\rm co}\mathcal{P}$ via a universal deterministic Turing machine $U^{\prime}$, because a polynomial-time deterministic Turing machine (we call it a $\mathcal{P}$ machine) that works under the “there exists no accepting path” condition is still a polynomial-time deterministic Turing machine. Namely, the “here exists at least one accepting path” condition and the “there exists no accepting path” condition for deterministic Turing machines are symmetrical. Specifically, for an arbitrary deterministic Turing machine $M$, if $M$ accepts the input $w$, then there is only one accepting path, and if $M$ rejects the input $w$, then there is also only one rejecting path. This is different from nondeterministic Turing machines. In short, if we apply the same technique, i.e., the simulation technique, to create a language not in ${\rm co}\mathcal{P}$, then the universal deterministic Turing machine $U^{\prime}$ (it runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$) needed to be constructed accepts the language $L(U^{\prime})\in\mathcal{P}$ (the proof is similar to the case in Section 5 below). But by the symmetrical properties between the “there exists at least one accepting path” condition and the “there exists no accepting path” condition with respect to deterministic Turing machines, we clearly have that $L(U^{\prime})\in{\rm co}\mathcal{P}$. Thus, we are unable to claim that $L(U^{\prime})\notin{\rm co}\mathcal{P}$.
@@ 6. Breaking the “Relativization Barrier”
We make the following specific assumptions:
@@ 7. Rich Structure of ${\rm co}\mathcal{NP}$
It is sufficient to show that for every language $L\in{\rm co}\mathcal{NP}$, $L\leq_{p}{\rm TAUT}$, where $\leq_{p}$ is polynomial-time many-one reduction [Kar72]. We can modify the Cook-Levin reduction [Coo71, Lev73] from $\overline{L}=\{0,1\}^{*}\setminus L$ to $\rm SAT$} (recall that all nondeterministic Turing machines are with input alphabet $\{0,1\}$ because that will be all we need; see Section 3). So for any input $w\in\{0,1\}^{*}$, the Cook-Levin reduction produces a formula $\varphi_{w}$ that is satisfiable if and only if $w\in\overline{L}$. In other words, the formula $\neg\varphi_{x}$ is in $\rm TAUT$} if and only if $x\in L$, which completes the proof.
@@
Then $L\in{\rm co}\mathcal{NP}$, and we will show that $L$ is not in $\mathcal{P}$ nor ${\rm co}\mathcal{NP}$-complete. To see so, suppose that $L\in\mathcal{P}$; then since $\mathcal{P}={\rm co}\mathcal{P}$, we have $\overline{L}=L_{\rm inter}\in\mathcal{P}$, a contradiction to the fact that $L_{\rm inter}\notin\mathcal{P}$ by Lemma 7.1. Further assume that $L$ is ${\rm co}\mathcal{NP}$-complete; then ${\rm TAUT}\leq_{p}L$. By Lemma 7.2, $\overline{L}=L_{\rm inter}$ is $\mathcal{NP}$-complete; it is also a contradiction to the fact that $L_{\rm inter}$ is not $\mathcal{NP}$-complete. Thereby, to summarize from the above two cases, we have shown that $L$ is ${\rm co}\mathcal{NP}$-intermediate. This finishes the proof.
@@ 9. Open Problems
In addition, does the complexity class $\mathcal{BQP}$ (see page 213 of [AB09]) contain the problem Satisfiability? Note that by Corollary 1.4, we have $\mathcal{BQP}\neq\mathcal{NP}$, but this cannot exclude that $\mathcal{NP}\subsetneqq\mathcal{BQP}$. So, if one is able to show that the problem Satisfiability is not in $\mathcal{BQP}$, then the precise relationships between $\mathcal{BQP}$ and $\mathcal{NP}$ would be clear. Currently, we have no additional information about this problem. We tried to think about it but also failed, and we remark that this is also a fascinating open problem in complexity theory.

### v23

@@ Abstract.
We prove in this paper that there is a language $L_{s}$ accepted by some nondeterministic Turing machine that runs within time $O(n^{k})$ for any positive integer $k\in\mathbb{N}_{1}$ but not by any ${\rm co}\mathcal{NP}$ machines. Then we further show that $L_{s}$ is in $\mathcal{NP}$, thus proving a groundbreaking result that
@@ 1.1. Our Contributions
$\mathcal{P}\neq\mathcal{NP}$. [^1] So, at the end of this paper, the reader will become clear that resolving the $\mathcal{P}\overset{?}{=}\mathcal{NP}$ problem does not necessarily require diagonalization techniques.
@@
Furthermore, Theorem 1.1 also implies that polynomial-time many-one reduction [Kar72] is more precise than polynomial-time Turing reduction [Coo71], since under the polynomial-time Turing reduction, every ${\rm co}\mathcal{NP}$ language is polynomial-time Turing reducible to an $\mathcal{NP}$ language, i.e., ${\rm co}\mathcal{NP}\subseteq\mathcal{NP}$. However, in fact, the complement of $L_{s}$ is in ${\rm co}\mathcal{NP}$ (i.e., $\overline{L_{s}}\in{\rm co}\mathcal{NP}$) but is not in $\mathcal{NP}$ (i.e., $\overline{L_{s}}\notin\mathcal{NP}$) by Theorem 1.1.
@@ 1.2. Background and Prior Work
In history, the fundamental measure of time opened the door to the study of the extremely expressive time complexity class $\mathcal{NP}$, one of the most important classical complexity classes, i.e., nondeterministic polynomial-time. Specifically, $\mathcal{NP}$ is the set of decision problems for which the problem instances, where the answer is “yes,” have proofs verifiable in polynomial time by some deterministic Turing machine, or alternatively the set of problems that can be solved in polynomial time by a nondeterministic Turing machine (see e.g., [A2]). The famous Cook-Levin theorem [Coo71, Lev73] shows that this class has complete problems, which states that Satisfiability is $\mathcal{NP}$-complete, i.e., Satisfiability is in $\mathcal{NP}$ and any other language in $\mathcal{NP}$ can be reduced to it in polynomial time. Meanwhile, it is also worth noting that this famous and seminal result stimulated the seminal and famous work [Kar72] of Karp, opening the door to the research of the renowned rich theory of $\mathcal{NP}$-completeness [Kar72].
@@
In the next few paragraphs, let us recall some of the interesting background, the current status, the history, and the main goals of the field of proof complexity.
@@
On the one hand, as we all know well, proof theory is a major branch of mathematical logic and theoretical computer science in which proofs are treated as formal mathematical objects, facilitating their analysis by mathematical techniques (see e.g., [A4]). Some of the major areas of proof theory include structural proof theory, ordinal analysis, automated theorem proving, and proof complexity, and so forth. It should be pointed out that much research also focuses on applications in computer science, linguistics, and philosophy, from which we see that the impact of proof theory is very profound. Moreover, as pointed out in [Raz04] by Razborov, proof and computations are among the most fundamental concepts relevant to virtually any human intellectual activity. Both have been central to the development of mathematics for thousands of years. The effort to study these concepts themselves in a rigorous, metamathematical way was initiated in the $20$th century and led to the flourishing of mathematical logic and derived disciplines.
@@
On the other hand, according to the points of view given in [Rec76], logicians have proposed a great number of systems for proving theorems, which give certain rules for constructing proofs and for associating a theorem (formula) with each proof. More importantly, these rules are much simpler to understand than the theorems. Thus, a proof gives a constructive way of understanding that a theorem is true. In addition, a proof system is sound if every theorem is true, and it is complete if every true statement (from a certain class) is a theorem (i.e., has a proof); see e.g., [Rec76].
@@
As we can see, the relationship between proof theory and proof complexity is inclusion. Within the area of proof theory, proof complexity is the subfield aiming to understand and analyze the computational resources that are required to prove or refute statements (see e.g., [A5]). Research in proof complexity is predominantly concerned with proving proof-length lower and upper bounds in various propositional proof systems. For example, among the major challenges of proof complexity is showing that the Frege system, the usual propositional calculus, does not admit polynomial-size proofs of all tautologies; see e.g., [A5]. Here the size of the proof is simply the number of symbols in it, and a proof is said to be of polynomial size if it is polynomial in the size of the tautology it proves. Moreover, contemporary proof complexity research draws ideas and methods from many areas in computational complexity, algorithms, and mathematics. Since many important algorithms and algorithmic techniques can be cast as proof search algorithms for certain proof systems, proving lower bounds on proof sizes in these systems implies run-time lower bounds on the corresponding algorithms; see e.g., [A5].
@@
Meanwhile, propositional proof complexity is an area of study that has seen a rapid development over the last two decades, which plays as important a role in the theory of feasible proofs as the role played by the complexity of Boolean circuits in the theory of efficient computations; see, e.g., the celebrated work [Raz15] by Razborov. In most cases, according to [Raz15], the basic question of propositional proof complexity can be reduced to that given a mathematical statement encoded as a propositional tautology $\psi$ and a class of admissible mathematical proofs formalized as a propositional proof system $P$, what is the minimal possible complexity of a $P$-proof of $\psi$? In other words, propositional proof complexity aims to understand and analyze the computational resources required to prove propositional tautologies, in the same way that circuit complexity studies the resources required to compute Boolean functions. In light of the point of view given in [Raz15], the task of proving lower bounds for strong proof systems like Frege or Extended Frege modulo any hardness assumption in the purely computational world may be almost as interesting; see e.g., [Raz15]. It is worth noting that it has been observed by Razborov [Raz15] that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$ implies lower bounds for any propositional proof system. In addition, see, e.g., [Raz03] for interesting directions in propositional proof complexity.
@@
Let us return back to the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem. As a matter of fact, the so-called $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem is the central problem in proof complexity; see, e.g., the excellent survey [Kra19]. It formalizes the question of whether or not there are efficient ways to prove the negative cases in ours and in many other similar examples. Moreover, the ultimate goal of proof complexity is to show that there is no universal propositional proof system allowing for efficient proofs of all tautologies, which is equivalent to showing that the computational complexity class $\mathcal{NP}$ is not closed under the complement; see e.g., [Kra19].
Furthermore, in 1974, Cook and Reckhow showed in [CR74] that there exists a super proof system if and only if $\mathcal{NP}$ is closed under complement; that is, if and only if $\mathcal{NP}={\rm co}\mathcal{NP}$. It is, of course, open whether any superproof system exists. The above result has led to what is sometimes called “Cook’s program” for proving $\mathcal{NP}\neq{\rm co}\mathcal{NP}$: prove superpolynomial lower bounds for proof lengths in stronger and stronger proposition proof systems until they are established for all abstract proof systems. Cook’s program, which also can be seen as a program to prove $\mathcal{P}\neq\mathcal{NP}$, is an attractive and plausible approach; unfortunately, it has turned out to be quite hard to establish superpolynomial lower bounds on common proof systems; see, e.g., the excellent survey on proof complexity [Bus12].
@@ 2. Preliminaries
There exists a Turing machine $U$ such that for every $x,\alpha\in\{0,1\}^{*}$, $U(x,\alpha)=M_{\alpha}(x)$, where $M_{\alpha}$ denotes the Turing machine represented by $\alpha$. Moreover, if $M_{\alpha}$ halts on input $x$ within $T(|x|)$ steps, then $U(x,\alpha)$ halts within $cT(|x|)\log T(|x|)$ steps, [^5] In this paper, $\log n$ stands for $\log_{2}n$. where $c$ is a constant independent of $|x|$ and depending only on $M_{\alpha}$’s alphabet size, number of tapes, and number of states.
@@ 3. All Polynomial-time Nondeterministic Turing Machines Are Enumerable
Figure 2. Cantor pairing function
There is another way to enumerate all polynomial-time nondeterministic Turing machines without encoding the order of polynomial into their representation. To do so, we need the Cantor pairing function (see Fig. 2 above from [A8]):
@@ 4. Simulation of All Polynomial-time Nondeterministic Turing Machines
One may wish to apply simulation techniques to the case of $\mathcal{P}$ versus ${\rm co}\mathcal{P}$ and doubt that simulation techniques are unable to produce a language not in ${\rm co}\mathcal{P}$. In this scenario, we remark that the simulation technique is unable to create a language not in ${\rm co}\mathcal{P}$ via a universal deterministic Turing machine $U^{\prime}$, because a polynomial-time deterministic Turing machine (we call it a $\mathcal{P}$ machine) that works under the “there exists no accepting path” condition is still a polynomial-time deterministic Turing machine. Namely, the “there exists at least one accepting path” condition and the “there exists no accepting path” condition for deterministic Turing machines are symmetrical. Specifically, for an arbitrary deterministic Turing machine $M$, if $M$ accepts the input $w$, then there is only one accepting path, and if $M$ rejects the input $w$, then there is also only one rejecting path. This is different from nondeterministic Turing machines. In short, if we apply the same technique, i.e., the simulation technique, to create a language not in ${\rm co}\mathcal{P}$, then the universal deterministic Turing machine $U^{\prime}$ (it runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$) needed to be constructed accepts the language $L(U^{\prime})\in\mathcal{P}$ (the proof is similar to the case in Section 5 below). But by the symmetrical properties between the “there exists at least one accepting path” condition and the “there exists no accepting path” condition with respect to deterministic Turing machines, we clearly have that $L(U^{\prime})\in{\rm co}\mathcal{P}$. Thus, we are unable to claim that $L(U^{\prime})\notin{\rm co}\mathcal{P}$.
@@ 6. Breaking the “Relativization Barrier”
Based on the above, we can make the following specific assumptions:
@@ 7. Rich Structure of ${\rm co}\mathcal{NP}$
It is sufficient to show that for every language $L\in{\rm co}\mathcal{NP}$, $L\leq_{m}^{p}{\rm TAUT}$, where $\leq_{m}^{p}$ is polynomial-time many-one reduction [Kar72]. We can modify the Cook-Levin reduction [Coo71, Lev73] from $\overline{L}=\{0,1\}^{*}\setminus L$ to $\rm SAT$} (recall that all nondeterministic Turing machines are with input alphabet $\{0,1\}$ because that will be all we need; see Section 3). So for any input $w\in\{0,1\}^{*}$, the Cook-Levin reduction produces a formula $\varphi_{w}$ that is satisfiable if and only if $w\in\overline{L}$. In other words, the formula $\neg\varphi_{x}$ is in $\rm TAUT$} if and only if $x\in L$, which completes the proof.
@@
Then $L\in{\rm co}\mathcal{NP}$, and we will show that $L$ is not in $\mathcal{P}$ nor ${\rm co}\mathcal{NP}$-complete. To see so, suppose that $L\in\mathcal{P}$; then since $\mathcal{P}={\rm co}\mathcal{P}$, we have $\overline{L}=L_{\rm inter}\in\mathcal{P}$, a contradiction to the fact that $L_{\rm inter}\notin\mathcal{P}$ by Lemma 7.1. Further assume that $L$ is ${\rm co}\mathcal{NP}$-complete; then ${\rm TAUT}\leq_{m}^{p}L$. By Lemma 7.2, $\overline{L}=L_{\rm inter}$ is $\mathcal{NP}$-complete; it is also a contradiction to the fact that $L_{\rm inter}$ is not $\mathcal{NP}$-complete. Thereby, to summarize from the above two cases, we have shown that $L$ is ${\rm co}\mathcal{NP}$-intermediate. This finishes the proof.
@@ 9. Open Problems
In addition, does the complexity class $\mathcal{BQP}$ (see page 213 of [AB09]) contain the problem Satisfiability (which requires someone to show that the problem Satisfiability can be solved by some bounded-error quantum polynomial-time Turing machines)? Note that by Corollary 1.4, we have $\mathcal{BQP}\neq\mathcal{NP}$, but this cannot exclude that $\mathcal{NP}\subsetneqq\mathcal{BQP}$. So, if one is able to show that the problem Satisfiability or the other problem in $\mathcal{NP}$ is not in $\mathcal{BQP}$, then the precise relationships between $\mathcal{BQP}$ and $\mathcal{NP}$ would be clear. [^9] It is widely believed that quantum polynomial-time Turing machines are unable to solve the problem Satisfiability with an error probability of at most $\frac{1}{3}$ for all instances. Currently, we have no additional information about this problem. We tried to think about it but also failed, and we remark that this is also a fascinating open problem in complexity theory.